OpenStudy (nincompoop):

find the areas of triangles enclosed by axes and tangent to y = 1/x

OpenStudy (nincompoop):

obtain the general slope of the tangent line for equation y = 1/x y' = (u'v - uv')/v^2 = (0*x0 - 1*1)/x0^2 -> y' = -1/x0^2 equation of the tangent line would then be: y-y0 = y'(x - x0) y -y0 = -x/x^2 )+ x0/x^2 locate the intercepts: y = 0 0 = -x/x^2 + x0/x^2 + 1/x0 x = 2x0

OpenStudy (nincompoop):

y = 2y0 Area = 1/2 (2x0 * 2y0) = 2 area = 2

OpenStudy (nincompoop):

yes, no, maybe so?

OpenStudy (abb0t):

No, \(\sf \color{red}{y'=-\frac{1}{x^2}}\)

OpenStudy (nincompoop):

x naut

OpenStudy (nincompoop):

-1/xnaut^2

OpenStudy (nincompoop):

laughing out loud

OpenStudy (nincompoop):

instead of assigning (a, f(a)) I am maintaining x's and y's

OpenStudy (abb0t):

I would switch to \(a\). Lol. In that case, you have \(\sf -\frac{1}{a^2}\) \(\Rightarrow\)y = \(\sf -\frac{1}{a^2}x+b\) which is your slope equation

OpenStudy (abb0t):

Now you can further proceed to solve for \(b\)

OpenStudy (abb0t):

y = \(\sf \frac{1}{x}\) and x = a or \(\sf x_0\), whatever u wanna call that sh*t.

OpenStudy (nincompoop):

laughing out loud yeah I did that but it my classmates had a hard time following

OpenStudy (abb0t):

after you get your \(b\), you can integrate

OpenStudy (nincompoop):

we don't have to integrate